OpenStudy (yeval76):

***HELP PLEASE*** Find the non-extraneous solutions of the square root of the quantity x plus 9 minus 5 equals quantity x plus 4. x = −2 x = −9 and x = −2 x = 9 x = −9 and x = −8

OpenStudy (anonymous):

\[\sqrt{x+9}-5=x+4 \] ?

OpenStudy (anonymous):

#4, x=-9 and x=-8

OpenStudy (yeval76):

Thank you!

OpenStudy (anonymous):

With a problem like this one, put the radical expressions on one side the rest on the other side. then square both sides to get rid of the radicals.

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (yeval76):

Can I ask you another question? @robtobey